# Toward Human-AI Alignment in Large-Scale Multi-Player Games

### v1

@@ Abstract
Machine Learning, ICML
Microsoft Research, $*{}^{*}$susharma@mit.edu
@@ 1 Introduction
Human-AI alignment is pivotal in generative AI research for several compelling reasons. First, as generative AI is increasingly integrated into various applications (Park et al., 2023; Brynjolfsson et al., 2023), ensuring alignment with human values and intentions becomes crucial to mitigate risks and enhance user trust (Sucholutsky et al., 2023; Gabriel, 2020). Second, aligning AI systems with human behavior fosters more effective collaboration between humans and machines (Chakraborti & Kambhampati, 2018), unlocking the potential for synergistic outcomes (Wynn et al., 2023; Bobu et al., 2023). Third, in ethical considerations surrounding the deployment of generative models, alignment serves as a safeguard against unintended consequences and biases, promoting responsible AI development and deployment (Kenthapadi et al., 2023; Weidinger et al., 2021). Thus, human-AI alignment in generative AI is essential for creating trustworthy, beneficial, and ethically sound AI applications that align with societal values and expectations.
The evaluation of human-agent alignment in rich observation and multi-agent environments poses a significant challenge (Leike et al., 2018; Ouyang et al., 2022; Wang et al., 2022, 2023b; Burns et al., 2023). In complex multi-agent video games, where each player faces a multitude of actions, this challenge involves finding the appropriate level of abstraction for a meaningful interpretation of human actions to evaluate artificial agents’ alignment.
In this work, we propose an interpretable approach towards human-AI alignment by introducing the “Task-sets” framework, offering a means to abstract task sets from the environment. Task-sets (Sakai, 2008) offer a higher level of abstraction compared to policies (Sutton et al., 1999a; Silver et al., 2014; Lillicrap et al., 2015; Schulman et al., 2017) and options (Sutton et al., 1999b; Precup, 2000; Stolle & Precup, 2002; Bacon et al., 2017; Khetarpal et al., 2020) in reinforcement learning.
@@
We use the task-set abstraction to interpret differences between agents as higher-level behavioral concepts that transcend comparing changes in policies and options alone. This abstraction also allows comparing behavior across temporal scales (Monsell, 2003; Sakai, 2008; Collins & Frank, 2013; Momennejad & Haynes, 2013; Vaidya & Badre, 2022). Task sets also facilitate compositionality. In the example above, the “walk to kitchen when hungry to find a snack” task-set could also be used to grab a snack for a visitor (taking their preferences into account) or to modify one’s own snack choice to account for healthiness. In Bleeding Edge, we use task-sets to abstract from actions and policies of players to higher-level notions, such as the dimensions of the behavior manifold (Fig.1a). This abstraction not only enables understanding of human cognitive processes, but also, fosters strong notions of transferability, both between agents and across environments, thus making it suitable for effective evaluation of alignment between AI agents and humans.
@@
Our research, driven by these investigative avenues, pursues two-fold primary objectives. 1) We seek a nuanced understanding of human cognition and behavior in the realm of large-scale multiplayer video games. 2) We aspire to harness this understanding to advance AI for gameplay, by constructing, evaluating and training artifical agents for targeted behavior replication through player style identification. In this work we mainly focus on the evaluation framework for measuring alignment (illustrated through a proof of concept AI agent). This framework can be used for evaluating alignment of any autonomous decision making AI agent (Yao et al., 2022; Sharma et al., 2022; Shinn et al., 2023; AutoGPT, 2023; Du et al., 2023; Wang et al., 2023a). In summary, we provide a framework for developing AI agents with superior alignment with humans.
@@ 2 Bleeding Edge
Gameplay: Bleeding Edge is designed for 4v4 multiplayer battles. This means that two teams, each consisting of four players, participate in each game. Players engage in team-based battles featuring dynamic combat dynamics, including both melee (close-quarters physical engagements) and ranged (attacks from a distance) elements. The multiplayer dynamics underscore the importance of teamwork, coordination, and communication in navigating the diverse and dynamic arenas in the game.
@@
Character selection and abilities: Players choose from a diverse roster of 13 characters, each with unique abilities and playstyles. Characters are classified into three categories (Fig.1c): Support, Tank, and Damage. Support characters focus on healing, buffs, and crowd control. Tanks absorb damage and protect teammates. Damage characters excel at eliminating opponents quickly. This allows for strategic team composition, promoting collaborative and strategic thinking (see appendix A.6 for details).
@@ 3 AI Agent
We trained an AI agent for playing Bleeding Edge and measured its alignment with humans. The model, with $\sim$ [^2] M parameters, is a transformer based architecture. We frame human gameplay trajectories as sequences of image-action pairs, and optimize the transformer to predict the next action in the sequence given the previous images. (Fig.2).
Observation Encoder: The model is trained on sequences of $T=128$ images where each image is reshaped to $128\times 128\times 3$ and then divided by 255 to ensure its value lies in range [0, 1]. A custom ResNet (He et al., 2016) with 18.6M parameters is used to embed each image observation ($O\in R^{3\times 128\times 128}$) independently into a vector (details in A.4.2). For each input image, the output of the encoder is a 1024D embedding.
Transformer: The causal transformer (with $\sim$203M parameters) is applied on the image embeddings ($Z\in R^{1024}$). Specifically, a GPT2-like architecture (Radford et al., 2019) from NanoGPT (Andrej Karpathy, 2023) was used containing 16 layers/blocks. Each attention layer has 8 heads with the action embeddings output of size 1024.
@@
Data and Training: We use Behavior cloning (Pomerleau, 1991) for training. For buttons we use the binary cross entropy loss with logits, and for the joysticks we use the cross entropy loss for each component. The total loss is computed as the sum of the losses for buttons and each of the joystick’s components. The training data was sampled from a dataset consisting of 57,661 full human gameplay videos where each video corresponds to continuous gameplay by one player, resulting in 1,707,997,180 video frames ($\sim 1.8$ billion time steps) and 7907.4 hours of human gameplay. The training took 6 days on 16 GPUs (V100s) and the model was trained for $\sim 72000$ steps (details in A.4.1).
AI Rollouts: We generated $600$ rollouts of $1$ min each with the above model by first picking a number of random game situations from the dataset. These are then filtered down to the desired characters e.g., Daemon, ZeroCool, and Makutu. For each game, we initialize the game at that state and let the AI agent play from that state onwards for a duration of $1$ minute. Except for the AI agent, all other players in the game are non-players who aren’t controlled and simply continue to repeat the latest action that the corresponding human player acted during the real human gameplay (e.g., if the player was moving forward when the game was originally recorded at that state, they will continue moving forward). They do not respond to the AI agent’s actions, however, their features get affected based on AI agent’s actions. For instance, their health decreases when the AI agent attacks them. We assume that each rollout is a different AI player given the stochasticity e.g., due to sampling the action from the output probability distribution.
@@ 4.1 Task-sets
Definition 4.1 (Task-Set).
A task-set comprises of extracting a set of features from the game state at each time step on which affordance and completion conditions are determined. Affordance conditions identify when a task-set can be performed or executed, giving the agent the choice of whether or not to engage in the task-set. Completion conditions determine if an agent successfully performed a task-set by choosing to engage. A task-set is said to be afforded when its affordance condition is met, and said to be completed with it’s completion condition is met.
@@
$Run\_From\_Enemy\_In\_Good\_Health$
@@
$Attack\_Approach\_Damage\_Enemy\_Health\_Good$
@@
We note that future work could consider using automatically learned task-sets similar to skill learning (Wang et al., 2023a; Khetarpal et al., 2021). However, in the scope of this work, we adhere to the programmer-specified definition of task sets based on our understanding of the game and the analysis of game play data, a deliberate choice aimed at illustrating the advantages inherent in this framework without being limited by the quality of learned task sets.
Cognitive themes: In our analytical framework, we employ three distinct cognitive themes that are ubiquitously observed in the behaviors of various biological species (Fig.3a). They are: 1)“Fight-Flight” for shedding light on the decision-making processes associated with confrontation and evasion, 2) “Explore-Exploit” is employed to discern the strategic balance between exploration and exploitation strategies within the game environment, and 3)“Solo-Multi-agent play” is used for understanding the interplay between individual and collaborative player behaviors. The task-sets defined for each of these cognitive themes are shown in Fig.3b (complete definitions in A.5).
@@ 4.2 Simultaneous affordance-completion analysis
    $P($completion at $t+x|$ simultaneous affordance at $t)$
@@ 5.1 Fight-Flight
For each pair of task-sets, we generate $9$ features: Area under the curve (AUC), Max, and Argmax of ‘fight’ task-set curve; AUC, Max, and Argmax of ‘flight’ task-set curve; and the ratio between each of the features, dividing fight / flight, resulting in $4\times 9=36$ interpretable features per player. We use these features to produce an unsupervised 2D embedding of each human player through UMAP (Uniform Manifold Approximation and Projection) (McInnes et al., 2018) as shown in Fig.4b, top. Here each point corresponds to one human player leading to a total of 123 points on the manifold. The data for each player is averaged across all the games (3 or more) played by this player. When the manifold is colored by the reward (score) received by the players, we find that the reward doesn’t shape the embedding space as shown in Fig.4b, bottom. However, when the manifold is colored by fight/flight AUC ratio, we find that the human behavioral manifold embedding space is shaped by fight or flight behavior as shown in Fig.4c. Although on an absolute scale, all players are flighty, relatively some players are more fighty (red points) than others.
@@
Conclusion: We find that while human players vary in how they express the fight or flight behavior while playing the same character, AI players do not.
[…]
@@ 5.2 Explore-Exploit
To test this hypothesis we defined three pairs of goal-directed navigation based task-sets listed in Fig.3b, middle and shown in Fig.5a. Exploitation prioritizes immediate rewards by efficiently achieving the objectives in a direct and goal-oriented manner. For instance, moving directly towards the nearest cluster to collect power cells, going directly to the nearest active platform for depositing them as quickly as possible, or going to the nearest inactive platform and waiting for it to become active.
Similar to the fight-flight analysis above (section 5.1), we compute the simultaneous affordance-completion curves for each pair of explore-exploit task-sets for human players as well as AI agents. Fig.5b,top shows that humans are exploiters - the exploit task-set has a higher completion probability than the explore task-set. Next, we compute the UMAP embedding as shown in Fig.5c,top. Again, we find that the reward doesn’t shape human behavioral manifold embedding space (Fig.5c, bottom), however, when colored by explore/exploit AUC ratio, the embedding space is shaped by explore or exploit behavior (Fig.5d). Although on an absolute scale, all players are exploiters, relatively some players explore more (red points) than others.
We ran the same analysis on the AI agent gameplay data. Fig. 5b,bottom shows the simultaneous affordance-completion curve showing that the AI has almost equal completion probability for the explore and exploit task-set. Fig. 4e,left shows the UMAP embedding of AI players, while playing Daemon. Fig.4e,middle shows the manifold colored by reward received, indicating that the reward doesn’t shape the AI embedding space. Fig.4e,right shows that when colored by explore/exploit AUC ratio, the AI embedding space is not shaped by explore-exploit behavior either. This indicates that although AI players show some diversity in exploring vs exploiting, both of these groups share similar feature representations, making it difficult to differentiate between them based on explore-exploit behavior. In other words, AI agents exhibit overlapping characteristics between exploration and exploitation behavior and can’t be grouped based on it. This is in contrast to the human players who are behaviorally different on a relative scale with some of them being more of explorers and some being more of exploiters (Fig.5d).
Conclusion: We find that while human players clearly vary in how they express explore or exploit behavior when playing the same character, we can’t differentiate between AI players based on this behavior. Moreover, the conclusion stands across the characters. [^1] We find that the same analysis for different characters in Fig.1c results in the same findings.
@@ 5.3 Solo-Multi-agent play
To identify differences in play-style across character types, we compute the overlap in task-set affordances and completions as shown in Fig.6a for each character type (Fig.A.8a, A.11a). On analysing the difference between these overlap matrices (Fig.A.8b, A.11b), we find that the Tank characters are the most suited to carrying power cells (Fig.A.10), Support characters to healing (Fig.A.12), whereas Damage characters to dealing damage (Fig.A.13).
From the completions overlap matrices, we extract the overlap of completions of solo-multi-agent game play task sets (Fig.3b, left) with the fight task-sets in the human gameplay data. We find that irrespective of the character type, all human players engage in fights more often in the presence of allies relative to when they are playing solo. Additionally, we find that the Damage characters are the most fighty.
Besides, we find that all the solo-mutli-agent task-sets are afforded simultaneously as is evident by the three identical affordance planes shown in Fig.6c,left. However, all of them have orthogonal completions illustrated by the three distinct completion planes (see Fig.6c, middle and right). This figure plots the data for the games played by the Daemon character. The table in the figure summarizes the fraction of times the three task-sets are afforded and completed (relative to affordances) in the game. Although all of them are simultaneously afforded, their completions are orthogonal, leading to completion fractions that sum up to $100$ for the three task sets. When playing the Daemon character, humans play in cooperation with their allies $\approx 56\%$ of the time, and play solo only $\approx 44\%$ of the time.

### v2

@@ Abstract
Toward Human-AI Alignment in Large-Scale Multi-Player Games
Sugandha Sharma1,2, Guy Davidson1,3, Khimya Khetarpal1,4, Anssi Kanervisto1, Udit Arora 1, Katja Hofmann1, Ida Momennejad1, 1Microsoft Research, 2Massachusetts Institute of Technology, 3New York University, 4Mila, University of Montreal Correspondence: susharma@mit.edu Published in Wordplay @ ACL 2024, Association for Computational Linguistics
@@ 1 Introduction
Human-AI alignment is pivotal in generative AI research for several compelling reasons. First, as generative AI is increasingly integrated into various applications Park et al. (2023); Brynjolfsson et al. (2023), ensuring alignment with human values and intentions becomes crucial to mitigate risks and enhance user trust Sucholutsky et al. (2023); Gabriel (2020). Second, aligning AI systems with human behavior fosters more effective collaboration between humans and machines Chakraborti and Kambhampati (2018), unlocking the potential for synergistic outcomes Wynn et al. (2023); Bobu et al. (2023). Third, in ethical considerations surrounding the deployment of generative models, alignment serves as a safeguard against unintended consequences and biases, promoting responsible AI development and deployment Kenthapadi et al. (2023); Weidinger et al. (2021). Thus, human-AI alignment in generative AI is essential for creating trustworthy, beneficial, and ethically sound AI applications that align with societal values and expectations.
The evaluation of human-agent alignment in rich observation and multi-agent environments poses a significant challenge Leike et al. (2018); Ouyang et al. (2022); Wang et al. (2022, 2023b); Burns et al. (2023). In complex multi-agent video games, where each player faces a multitude of actions, this challenge involves finding the appropriate level of abstraction for a meaningful interpretation of human actions to evaluate artificial agents’ alignment.
In this work, we propose an interpretable approach towards human-AI alignment by introducing the “Task-sets" framework, offering a means to abstract task sets from the environment. Task-sets Sakai (2008) offer a higher level of abstraction compared to policies Sutton et al. (1999a); Silver et al. (2014); Lillicrap et al. (2015); Schulman et al. (2017) and options Sutton et al. (1999b); Precup (2000); Stolle and Precup (2002); Bacon et al. (2017); Khetarpal et al. (2020) in reinforcement learning.
@@
We use the task-set abstraction to interpret differences between agents as higher-level behavioral concepts that transcend comparing changes in policies and options alone. This abstraction also allows comparing behavior across temporal scales Monsell (2003); Sakai (2008); Collins and Frank (2013); Momennejad and Haynes (2013); Vaidya and Badre (2022). Task sets also facilitate compositionality. In the example above, the “walk to kitchen when hungry to find a snack” task-set could also be used to grab a snack for a visitor (taking their preferences into account) or to modify one’s own snack choice to account for healthiness. In Bleeding Edge, we use task-sets to abstract from actions and policies of players to higher-level notions, such as the dimensions of the behavior manifold (Fig.1a). This abstraction not only enables understanding of human cognitive processes, but also, fosters strong notions of transferability, both between agents and across environments, thus making it suitable for effective evaluation of alignment between AI agents and humans.
@@
Our research, driven by these investigative avenues, pursues two-fold primary objectives. 1) We seek a nuanced understanding of human cognition and behavior in the realm of large-scale multiplayer video games. 2) We aspire to harness this understanding to advance AI for gameplay, by constructing, evaluating and training artifical agents for targeted behavior replication through player style identification. In this work we mainly focus on the evaluation framework for measuring alignment (illustrated through a proof of concept AI agent). This framework can be used for evaluating alignment of any autonomous decision making AI agent Yao et al. (2022); Sharma et al. (2022); Shinn et al. (2023); AutoGPT (2023); Du et al. (2023); Wang et al. (2023a). In summary, we provide a framework to evaluate human-AI alignment that could potentially be applied towards developing AI agents with superior alignment with humans.
@@ 2 Bleeding Edge
Gameplay: Bleeding Edge is designed for 4v4 multiplayer battles. This means that two teams, each consisting of four players, participate in each game. Players engage in team-based battles featuring dynamic combat dynamics, including both melee (close-quarters physical engagements) and ranged (attacks from a distance) elements.
@@
Character selection and abilities: Players choose from a diverse roster of 13 characters, each with unique abilities and playstyles. Characters are classified into three categories (Fig.1c): Support, Tank, and Damage (see appendix A.5 for details).
@@ 3 AI Agent
We trained an AI agent for playing Bleeding Edge and measured its alignment with humans. The model, with $\sim$ [^2] M parameters, is a transformer based architecture. We frame human gameplay trajectories as sequences of image-action pairs, and optimize the transformer to predict the next action in the sequence given the previous images (Fig.2).
Observation Encoder: The model is trained on sequences of $T=128$ images where each image is reshaped to $128\times 128\times 3$ and then divided by 255 to ensure its value lies in range [0, 1]. A custom ResNet He et al. (2016) with 18.6M parameters is used to embed each image observation ($O\in R^{3\times 128\times 128}$) independently into a vector (details in A.3.2). For each input image, the output of the encoder is a 1024D embedding.
Transformer: The causal transformer (with $\sim$203M parameters) is applied on the image embeddings ($Z\in R^{1024}$). Specifically, a GPT2-like architecture Radford et al. (2019) from NanoGPT Andrej Karpathy (2023) was used containing 16 layers/blocks. Each attention layer has 8 heads with the action embeddings output of size 1024.
@@
Data and Training: We use Behavior cloning Pomerleau (1991) for training. For buttons we use the binary cross entropy loss with logits, and for the joysticks we use the cross entropy loss for each component. The total loss is computed as the sum of the losses for buttons and each of the joystick’s components. The training data was sampled from a dataset consisting of 57,661 full human gameplay videos where each video corresponds to continuous gameplay by one player, resulting in 1,707,997,180 video frames ($\sim 1.8$ billion time steps) and 7907.4 hours of human gameplay. The training took 6 days on 16 GPUs (V100s) and the model was trained for $\sim 72000$ steps (details in A.3.1).
AI Rollouts: We generated $600$ rollouts of $1$ min each with the above model by first picking a number of random game situations from the dataset. These are then filtered down to the desired characters e.g., Daemon, ZeroCool, and Makutu. (see appendix A.6 for details).
@@ 4.1 Task-sets
Definition 1 (Task-Set).
A task-set comprises of extracting a set of features from the game state at each time step on which affordance and completion conditions are determined. Affordance conditions identify when a task-set can be performed or executed, giving the agent the choice of whether or not to engage in the task-set. Completion conditions determine if an agent successfully performed a task-set by choosing to engage. A task-set is said to be afforded when its affordance condition is met, and said to be completed when it’s completion condition is met.
@@
Run_From_Enemy_In_Good_Health
@@
Attack_Approach_Damage_Enemy_Health_Good
@@
We note that future work could consider using automatically learned task-sets similar to skill learning Wang et al. (2023a); Khetarpal et al. (2021). However, in the scope of this work, we adhere to the programmer-specified definition of task sets based on our understanding of the game and the analysis of game play data, a deliberate choice aimed at illustrating the advantages inherent in this framework without being limited by the quality of learned task sets.
Cognitive themes: In our analytical framework, we employ three distinct cognitive themes that are ubiquitously observed in the behaviors of various biological species (Fig.3a). They are: 1)“Fight-Flight" for shedding light on the decision-making processes associated with confrontation and evasion, 2) “Explore-Exploit" is employed to discern the strategic balance between exploration and exploitation strategies within the game environment, and 3)“Solo-Multi-agent play" is used for understanding the interplay between individual and collaborative player behaviors. The task-sets defined for each of these cognitive themes are shown in Fig.3b (complete definitions in A.4).
@@ 4.2 Simultaneous affordance-completion analysis
$P($completion at $t+x|$ simul. afford. at $t)$
@@ 5.1 Fight-Flight
For each pair of task-sets, we generate $9$ features: Area under the curve (AUC), Max, and Argmax of ‘fight’ task-set curve; AUC, Max, and Argmax of ‘flight’ task-set curve; and the ratio between each of the features, dividing fight / flight, resulting in $4\times 9=36$ interpretable features per player. We use these features to produce an unsupervised 2D embedding of each human player through UMAP (Uniform Manifold Approximation and Projection) McInnes et al. (2018) as shown in Fig.4b, top. Here each point corresponds to one human player leading to a total of 123 points on the manifold. The data for each player is averaged across all the games (3 or more) played by this player. When the manifold is colored by the reward (score) received by the players, we find that the reward doesn’t shape the embedding space as shown in Fig.4b, bottom. However, when the manifold is colored by fight/flight AUC ratio, we find that the human behavioral manifold embedding space is shaped by fight or flight behavior as shown in Fig.4c. Although on an absolute scale, all players are flighty, relatively some players are more fighty (red points) than others.
@@
Conclusion: We find that while human players vary in how they express the fight or flight behavior while playing the same character, AI players do not. Apart from Daemon, we also ran the above analysis for other characters in Fig.1c and drew the same conclusion irrespective of the character.
@@ 5.2 Explore-Exploit
To test this hypothesis we defined three pairs of goal-directed navigation based task-sets listed in Fig.3b, middle and shown in Fig.5a. Exploitation prioritizes immediate rewards by efficiently achieving the objectives in a direct and goal-oriented manner. Similar to the fight-flight analysis above (section 5.1), we compute the simultaneous affordance-completion curves for each pair of explore-exploit task-sets for human players as well as AI agents. Fig.5b,top shows that humans are exploiters - the exploit task-set has a higher completion probability than the explore task-set. Next, we compute the UMAP embedding as shown in Fig.5c,top. Again, we find that the reward doesn’t shape human behavioral manifold embedding space (Fig.5c, bottom), however, when colored by explore/exploit AUC ratio, the embedding space is shaped by explore or exploit behavior (Fig.5d). Although on an absolute scale, all players are exploiters, relatively some players explore more (red points) than others.
We ran the same analysis on the AI agent gameplay data. Fig. 5b,bottom shows that the AI has almost equal completion probability for the explore and exploit task-set. Fig. 4e,left shows the UMAP embedding of AI players, while playing Daemon. Fig.4e,middle shows the manifold colored by reward received, indicating that the reward doesn’t shape the AI embedding space. Fig.4e,right shows that when colored by explore/exploit AUC ratio, the AI embedding space is not shaped by explore-exploit behavior either. This indicates that although AI players show some diversity in exploring vs exploiting, both of these groups share similar feature representations, making it difficult to differentiate between them based on explore-exploit behavior. In other words, AI agents exhibit overlapping characteristics between exploration and exploitation behavior and can’t be grouped based on it. This is in contrast to the human players who are behaviorally different on a relative scale with some of them being more of explorers and some being more of exploiters (Fig.5d).
Conclusion: We find that while human players clearly vary in how they express explore or exploit behavior when playing the same character, we can’t differentiate between AI players based on this behavior. This conclusion stands across characters. [^1] We find that the same analysis for different characters in Fig.1c results in the same findings.
@@ 5.3 Solo-Multi-agent play
To identify differences in play-style across character types, we compute the overlap in task-set affordances and completions (Fig.6a) for each character type (Fig.A.8a, A.11a). On analysing the difference between these overlap matrices (Fig.A.8b, A.11b), we find that the Tank characters are the most suited to carrying power cells (Fig.A.10), Support characters to healing (Fig.A.12), and Damage characters to dealing damage (Fig.A.13). From the completions overlap matrices, we extract the overlap of completions of solo-multi-agent game play task sets (Fig.6b, left) with the fight task-sets in the human gameplay data. We find that irrespective of the character type, all human players engage in fights more often in the presence of allies relative to when they are playing solo. Additionally, we find that the Damage characters are the most fighty.
Besides, we find that all the solo-mutli-agent task-sets are afforded simultaneously as is evident by the three identical affordance planes shown in Fig.6c,left. However, all of them have orthogonal completions illustrated by the three distinct completion planes (see Fig.6c, middle and right). The table in the figure summarizes the fraction of times the three task-sets are afforded and completed (relative to affordances) in the game. Although all of them are simultaneously afforded, their completions are orthogonal, leading to completion fractions that sum up to $100$ for the three task sets. When playing the Daemon character, humans play in cooperation with their allies $\approx 56\%$ of the time, and play solo only $\approx 44\%$ of the time.